# Aggregating Local and Global Features via Selective State Spaces Model for Efficient Image Deblurring

### v1

@@ IV-A1 Datasets
In this section, we will introduce the datasets used and provide details regarding the training configurations.
Image Motion Deblurring. Following recent methods [26, 6], we train ALGNet using the GoPro dataset [20], which includes 2,103 image pairs for training and 1,111 pairs for evaluation. To assess the generalizability of our approach, we directly apply the GoPro-trained model to the test images of the HIDE [43] and RealBlur [44] datasets. The HIDE dataset contains 2,025 images that collected for human-aware motion deblurring. Both the GoPro and HIDE datasets are synthetically generated, but the RealBlur dataset comprises image pairs captured under real-world conditions. This dataset includes two subsets: RealBlur-J, formed with camera JPEG outputs, and RealBlur-R, generated offline through processes like white balance adjustment, demosaicking, and denoising applied to RAW images.
@@ IV-A2 Training details
For different tasks, we train separate models, and unless specified otherwise, the following parameters are employed. The models are trained using the Adam optimizer [46] with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The initial learning rate is set to $5\times 10^{-4}$ and gradually reduced to $1\times 10^{-7}$ using the cosine annealing strategy [47]. The batch size is chosen as $32$, and patches of size $256\times 256$ are extracted from training images. Data augmentation involves horizontal and vertical flips. We scale the network width by setting the number of channels to 32 and 64 for ALGNet and ALGNet-B, respectively.
@@ IV-B1 Image Motion Deblurring
The quantitative results on the GoPro and HIDE dataset are shown in Table I II. Our ALGNet-B shows 0.43 dB performance improvement over NAFNet-64 [4] on GoPro [20]. Moreover, as shown in Figure 1, through the scaling up of the model size, our ALGNet achieves even better performance, highlighting the scalability of ALGNet. Noted that, even though our network is trained solely on the GoPro [20] dataset, it still achieves a substantial gain of 0.19 dB PSNR over Restormer-Local [16] on the HIDE [43] dataset. This demonstrates its impressive generalization capability. Figure. 4 shows our model recovered more visually pleasant results. We also evaluate our ALGNet on the real-world images of RealBlur dataset [44] and achieve the best performance as shown in Table III.
Figure 4: Image motion deblurring comparisons on the RealBlur dataset [44].
TABLE III: Quantitative real-world deblurring results.
@@
DeblurGAN-v2 [13] 36.44 0.935 29.69 0.870
MPRNet [6] 39.31 0.972 31.76 0.922
Stripformer [17] 39.84 0.975 32.48 0.929
MRLRFNet [19] 40.92 0.975 33.19 0.936
FFTformer [14] 40.11 0.973 32.62 0.932
ALGNet(ours) 41.21 0.977 33.38 0.937
TABLE IV: Quantitative comparisons with previous leading single-image defocus deblurring methods on the DPDD testset [45] (containing 37 indoor and 39 outdoor scenes).
@@ IV-B2 Single-Image Defocus Deblurring
Table IV presents a numerical comparison of defocus deblurring methods on the DPDD dataset [45]. ALGNet outperforms other state-of-the-art methods across most metrics. Notably, in the outdoor scenes category, ALGNet exhibits a 0.15 dB improvement over the leading method IRNeXt [3]. The visual results in Figure 5 illustrate that our method recovers more details and visually aligns more closely with the ground truth compared to other algorithms.
@@ IV-C Ablation Studies
Here we present ablation experiments to verify the effectiveness and scalability of our method. Evaluation is performed on the GoPro dataset [20], and the results are shown in Table. V. The baseline is NAFNet [4]. We perform the break-down ablation by applying the proposed modules to the baseline successively, we can make the following observations: a) The Local block and Global block do not lead to significant performance gains when used alone. This is because our baseline has been able to capture the local information well, so the local block cannot improve the representation ability of the model too much. However, when we replace the module in baseline with our proposed global block, due to the problems of pixel forgetting and channel redundancy, it cannot improve the performance significantly. b) When we use both Global Block and local Block, we can see a big improvement in performance, up 0.52dB compared to baseline. c) The FA makes a gain of 0.14 dB for our model. e) When all three of them are used together, the lifting capacity of each module is higher.
In addition, to examine the advantage of our FA design, we compare our FA with other method, such as sum and concatenation. As shown in Table VI, our FA achieves better results, which indicates that our FA is able to emphasize the importance of the local block in restoration, enabling our model to select the most detail information for recovery. It is worth mentioning that our design is extremely lightweight, as it does not introduce additional convolution layers.
TABLE V: Ablation study on individual components of the proposed ALGNet.
Global Block Local Block FA PSNR
[…]
TABLE VI: The impact of skip connections on the overall performance.
@@ IV-D Resource Efficient
We examine the model complexity of the proposed approach and the state-of-the-art methods in terms of model running time, and FLOPs. Table VII and Figure 1 demonstrate that our ALGNet model attains state-of-the-art performance while concurrently reducing computational costs. Specifically, on the GoPro dataset [20], we achieve a 0.2 dB improvement over the previous best approach, FSNet [26], while using with up to 84.7% of cost reduction and achieving nearly 1.5$\times$ faster inference. This highlights the efficiency of our method, showcasing superior performance along with resource effectiveness.
TABLE VII: The evaluation of model computational complexity on the GoPro dataset [20]. The FLOPs are evaluated on image patches with the size of 256×256 pixels. The running time is evaluated on images with the size of 1280 × 720 pixels.
@@ V Conclusion
In this paper,we propose an efficient image deblurring network that leverages selective structured state spaces model to aggregate enriched and accurate features. We design an aggregate local and global block (ALGBlock) consisting of both global and local components. The global block utilizes selective structured state spaces model to capture long-range dependency information. And the local block models local connectivity using simplified channel attention and address the issues of local pixel forgetting and channel redundancy in the state space equation. In addition, we present an aggregate methods that emphasize the importance of the local block in restoration. Experimental results demonstrate that the proposed method outperforms state-of-the-art approaches on widely used benchmarks.

### v2

@@ IV-A1 Datasets
Image Motion Deblurring. Following recent methods [28, 6], we train ALGNet using the GoPro dataset [20], which includes 2,103 image pairs for training and 1,111 pairs for evaluation. To assess the generalizability of our approach, we directly apply the GoPro-trained model to the test images of the HIDE [41] and RealBlur [42] datasets. The HIDE dataset contains 2,025 images that collected for human-aware motion deblurring. Both the GoPro and HIDE datasets are synthetically generated, but the RealBlur dataset comprises image pairs captured under real-world conditions. This dataset includes two subsets: RealBlur-J, and RealBlur-R.
@@ IV-A2 Training details
For various tasks, separate models are trained, and unless otherwise specified, the following parameters are utilized. The models are trained using the Adam optimizer [44] with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The initial learning rate is set to $5\times 10^{-4}$ and gradually reduced to $1\times 10^{-7}$ using the cosine annealing strategy [45]. The batch size is chosen as $32$, and patches of size $256\times 256$ are extracted from training images. Data augmentation involves horizontal and vertical flips. We scale the network width by setting the number of channels to 32 and 64 for ALGNet and ALGNet-B, respectively.
TABLE I: Quantitative evaluations of the proposed approach against state-of-the-art motion deblrrring methods. The best and second best scores are highlighted and underlined. Our ALGNet-B and ALGNet are trained only on the GoPro dataset.
GoPro [20] HIDE [41]
Methods PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
DeblurGAN-v2 [13] 29.55 0.934 26.61 0.875
MPRNet [6] 32.66 0.959 30.96 0.939
MPRNet-local [6] 33.31 0.964 31.19 0.945
HINet [7] 32.71 0.959 30.32 0.932
HINet-local [7] 33.08 0.962 - -
Uformer [18] 32.97 0.967 30.83 0.952
MSFS-Net [46] 32.73 0.959 31.05 0.941
MSFS-Net-local [46] 33.46 0.964 31.30 0.943
NAFNet-32 [4] 32.83 0.960 - -
NAFNet-64 [4] 33.62 0.967 - -
Restormer [16] 32.92 0.961 31.22 0.942
Restormer-local [16] 33.57 0.966 31.49 0.945
IRNeXt [3] 33.16 0.962 - -
SFNet [27] 33.27 0.963 31.10 0.941
FSNet [28] 33.29 0.963 31.05 0.941
DeblurDiNAT-S [47] 32.85 0.961 30.65 0.936
DeblurDiNAT-L [47] 33.42 0.965 31.28 0.943
MambaIR [24] 33.21 0.962 31.01 0.939
ALGNet(Ours) 33.49 0.964 31.64 0.947
ALGNet-B(Ours) 34.05 0.969 31.68 0.952
@@ IV-B1 Image Motion Deblurring
We present the performance of evaluated image deblurring approaches on the synthetic GoPro [20] and HIDE [41] datasets in Tables I. Our ALGNet-B demonstrates a 0.43 dB improvement in performance over NAFNet-64 [4] on the GoPro [20] dataset. Compared with MambaIR [24], which is also based on the state space model, our ALGNet demonstrates an improvement in performance by 0.28 dB, while ALGNet-B achieves a substantial improvement of 0.84 dB. Additionally, as depicted in Figure 1, our ALGNet achieves even better performance through the scaling up of the model size, underscoring the scalability of ALGNet. Despite being trained solely on the GoPro [20] dataset, our network still achieves a significant gain of 0.19 dB PSNR over Restormer-Local [16] on the HIDE [41] dataset, demonstrating its generalization capability. Figure 4 illustrates that our model produces visually more pleasing results.
We also evaluate our ALGNet on real-world images from the RealBlur dataset [42] under two experimental settings: (1) applying the GoPro-trained model directly on RealBlur, and (2) training and testing on RealBlur data. As shown in Table II, for setting 1, our ALGNet achieves performance gains of 0.16 dB on the RealBlur-R subset over Restormer [16] and 0.13 dB on the RealBlur-J subset over DeBlurDiNAT-L [47]. Compared with MambaIR [24], our gains are 0.37 dB and 0.30 dB on RealBlur-R and RealBlur-J, respectively. A similar trend is observed for setting 2, where our gains over MRLPFNet [19] are 0.24 dB on RealBlur-R. Although our ALGNet performs slightly inferiorly to MRLPNet in PSNR metric on the RealBlur-J dataset, our SSIM metric is higher. Moreover, for setting 1, our method outperforms MRLPNet, indicating superior generalization capability. Figure 5 presents visual comparisons of the evaluated approaches. Overall, the images restored by our model exhibit sharper details and are closer to the ground truth compared to those produced by other methods.
TABLE II: Quantitative real-world deblurring results under two different settings: 1). applying our GoPro trained model directly on the RealBlur dataset [42], 2). Training and testing on RealBlur data where methods are denoted with symbol***
@@
MPRNet [6] 35.99 0.952 28.70 0.873
Restormer [16] 36.19 0.957 28.96 0.879
Stripformer [17] 36.08 0.954 28.82 0.876
FFTformer [14] 35.87 0.953 27.75 0.853
MRLPFNet [19] 36.16 0.955 28.98 0.861
DeblurDiNAT-S [47] 35.92 0.954 28.80 0.877
DeblurDiNAT-L [47] 36.07 0.956 28.99 0.885
MambaIR [24] 35.98 0.955 28.82 0.875
ALGNet(Ours) 36.35 0.961 29.12 0.886
$*{}^{*}$ [13] 36.44 0.935 29.69 0.870
$*{}^{*}$ [6] 39.31 0.972 31.76 0.922
$*{}^{*}$ [17] 39.84 0.975 32.48 0.929
$*{}^{*}$ [14] 40.11 0.973 32.62 0.932
$*{}^{*}$ [19] 40.92 0.975 33.19 0.936
$*{}^{*}$ [24] 39.92 0.972 32.44 0.928
$*{}^{*}$(Ours) 41.16 0.981 32.94 0.946
TABLE III: Quantitative comparisons with other single-image defocus deblurring methods on the DPDD testset [43] (containing 37 indoor and 39 outdoor scenes).
@@
Figure 5: Image motion deblurring comparisons on the RealBlur dataset [42]. Our ALGNet recovers image with clearer details.
@@ IV-B2 Single-Image Defocus Deblurring
We conduct single-image defocus deblurring experiments on the DPDD [43] dataset. Table III presents image fidelity scores of state-of-the-art defocus deblurring methods. ALGNet outperforms other state-of-the-art methods across all scene categories. Notably, in the combined scenes category, ALGNet exhibits a 0.15 dB improvement over the leading method IRNeXt [3]. In comparison to MambaIR [24], which also relies on the state space model, our ALGNet showcases an improvement in performance by 0.48 dB in indoor scenes. The visual results in Figure 6 illustrate that our method recovers more details and visually aligns more closely with the ground truth compared to other algorithms.
@@ IV-C Ablation Studies
Figure 7: The outputs of the CLGF module are processed by ReLU and global average pooling to obtain channel activation values.
Here we present ablation experiments to verify the effectiveness and scalability of our method. Evaluation is performed on the GoPro dataset [20], and the results are shown in Table. IV. The baseline is NAFNet [4]. We perform the break-down ablation by applying the proposed modules to the baseline successively, we can make the following observations:
  1. 1.
    When our CLGF module consists of only one local branch or global branch, the improvement in deblurring performance is not significant. There are two main reasons for this. Firstly, our baseline model is already capable of fully capturing local information, rendering the addition of a local branch ineffective in enhancing the model’s representation ability. Secondly, since our global branch is based on the state-space model, although it can capture long-distance information, it often encounters issues such as local pixel loss and channel redundancy, as illustrated in Figure 3. Therefore, when used alone, it fails to enhance performance.
  2. 2.
    When our CLGF module comprises both a local branch and a global branch, we observe a significant improvement in performance, up to 0.52 dB compared to the baseline. This indicates that CLGF has the ability to capture long-range dependency features and model local connectivity effectively.
  3. 3.
    The FA contributes a gain of 0.14 dB to our model.
To further validate the effectiveness of our CLGF module, we apply ReLU and global average pooling operations on the output results of CLGF to obtain channel activation values (see Figure 7). It’s evident that our CLGF successfully circumvents the issue of channel redundancy caused by an excessive number of hidden states in the state spaces model.
Furthermore, to assess the advantage of our FA design, we compare it with other methods such as sum and concatenation. As shown in Table V, our FA consistently delivers superior results, indicating its effectiveness in emphasizing the importance of the local branch in restoration. Importantly, our design does not introduce any additional computational burden.
Finally, we compare the feature maps before and after our ALGBLock in Figure 8. It is evident that the feature map from our local branch contains more detailed information compared to that from the global branch. Upon aggregation of the local and global branches of CLGF using FA, we observe a significant recovery of more details, particularly in the blurred license plate number present in the initial feature map.
TABLE IV: Ablation study on individual components of the proposed ALGNet.
Method PSNR
Baseline 32.83
Baseline + CLGF w/o local branch 32.89
Baseline + CLGF w/o global branch 32.86
Baseline + CLGF 33.35
Baseline + CLGF + FA 33.49
TABLE V: The impact of feature aggregation method on the overall performance.
@@
Figure 8: The internal features of ALGBlock. With our CLGF and FA, ALGBlcok produces more fine details than the initial feature, e.g., the number plate. Zoom in for the best view..
@@ IV-D Resource Efficient
We assess the model complexity of our proposed approach and state-of-the-art methods in terms of model running time and FLOPs. Table VI and Figure 1 illustrate that our ALGNet model achieves SOTA performance while simultaneously reducing computational costs. Specifically, we achieve a 0.2 dB improvement over the previous best approach, FSNet [28], with up to 84.7% cost reduction and nearly 1.5 times faster inference. Compared to MambaIR [24], our ALGNet reduces computational costs by 96.1% and achieves 3.1 times faster inference. This underscores the efficiency of our method, demonstrating superior performance along with resource effectiveness.
TABLE VI: The evaluation of model computational complexity on the GoPro dataset [20]. The FLOPs are evaluated on image patches with the size of 256×256 pixels. The running time is evaluated on images with the size of 1280 × 720 pixels.
@@ V Conclusion
In this paper,we propose an efficient image deblurring network that leverages selective state spaces model to aggregate enriched and accurate features. We design an ALGBlock consisting of CLGF and FA module. The CLGF module captures long-range dependency features using a selective state spaces model in the global branch, while employing simplified channel attention to model local connectivity in the local branch, thus reducing local pixel forgetting and channel redundancy. Additionally, we propose the FA module to emphasize the significance of the local information by dynamically recalibrating the weights through a learnable factor during the aggregation of the CLGF two branches. Experimental results demonstrate that the proposed method outperforms state-of-the-art approaches.